# Acquiring Better Load Estimates by Combining Anomaly and Change-point Detection in Power Grid Time-series Measurements

### v1

@@ 1 Introduction
Next to expansion, smarter use of the existing grid is essential. This can be done through for example flexible energy contracts, the use of redundant grid capacity for power generation, and the use of batteries Sufyan et al. (2019). Better insight into grid usage over time is needed to facilitate these changes. With more measurements this can be achieved, but this means also more data that needs to be cleaned before it can be used for analysis.
In order to determine key points of expansion, an accurate overview of the current state of grid capacity needs to be made. In this study, we specifically study primary substation-level measurements on the Dutch power grid managed by Alliander. In order to know what percentage of a primary substation’s capacity is being used, accurate estimates of the minimum and maximum load of the subgrid that substation supplies must be made. This process is more commonly known as demand modelling (Betge et al., 2021). Primary substation-level time series load measurements, however, cannot be directly interpreted without giving a distorted view. These measurements are contaminated with anomalies, for example caused by measurement errors, and with switch events. In a switch event power from different primary substations is rerouted to or from part of the measured primary substation’s route. An example of this can be found in Figure 1, where due to a cable failure power is rerouted from primary substation 1 to supply secondary substations E and F. Rerouting power reduces the blackout time in case of damage to a cable in the power grid, and is possible when the grid is sufficiently redundant to circumvent the broken cable. Switch events can also happen when long-term maintenance is performed on a primary substation or any its cables or secondary substations. Switch events can have a large range of possible lengths, from a few minutes, to multiple months, depending on how fast the underlying problem is resolved. Both of these, measurement errors, which we’ll call anomalies, and switch events, need to be filtered out to get an accurate estimate of the load profile. This more accurate load profile can be used directly for more optimized usage of the grid. It can then be used to find the true minimum and maximum load of the substation under normal operating conditions, so the redundancy capacity can be added separately. The task of detecting which parts of a load measurement represent the normal situation can be seen as a time-series segmentation task.
@@
Traditionally, these load measurement are manually segmented and annotated by domain experts within the Alliander organisation. This procedure is however extremely time consuming, and produces annotations with a high percentage of label noise. This label noise is of little concern in determining the minimum and maximum loads, as the annotation procedure is generally optimized in such a way that these loads are as accurate as possible. The resulting annotation can however not be used for other purposes such as year-round usage insights. An automated annotation procedure, capable of producing high-quality annotations, can therefore both free up valuable domain expert time, as well as allow for detailed studies on load measurement data previously unfeasible. Because economically and logistically vital decisions are made based on these load estimates, it is extremely important that any machine learning model automating this procedure is highly interpretable.
For smart grids, change point and anomaly detection algorithms have been broadly applied. Zhang et al. (2021) describe multiple time-series anomaly detection algorithms used in various applications across smart grids. They differentiate between point, contextual, and pattern anomalies. Thomas et al. (2020) describe a novel method tailored to detecting islanding. In addition to islanding, their method, which uses $k$-means clustering and empirical mode decomposition, is able to detect load switch events and other faults. Neagoe et al. (2021) used change point detection in order to detect multi-year patterns in hydropower generation in Romania. Wang and Ahn (2020) combine a regression-based anomaly detection model with SVM, $k$NN and a cross-entropy loss function to detect anomalies in a Tanzanian solar power plant. Wang et al. (2023) present methodology for simultaneously performing load forecasting and semi-supervised anomaly detection and apply it to the UCI electric load dataset. They report higher specificity and sensitivity than neural network-based methods, but find similar F1 scores. For shorter time series, Rajabi et al. (2020) have compared various clustering methods for load pattern segmentation.
@@ 2.1 Data
In this study we optimize and evaluate our algorithms for anomaly and switch event detection on a total of $180$ primary substation load measurements. Most primary substation measurements span a full year in length, and provide measurements at regular 15-minute intervals of the apparent power $S$, which we’ll refer to as load. We calculate $S$ from the active power $P$ and the reactive power $Q$ and assign it the sign of $P$, $S=sign(P)*\sqrt{P^{2}+Q^{2}}$. In some cases, the measurement equipment of a primary substation does not allow for accurate measurements of both $P$ and $Q$. In these cases, $S$ is calculated from $S=\sqrt{3}*V*I$. The $\sqrt{3}$ term originates from the fact that in a 3-phase system the phase voltage rather than the line voltage is used.
@@
At each 15-minute interval over which the load is measured, we also calculate the so-called bottom-up load $B$ throughout the subgrid supplied by that specific primary substation. The bottom-up load is an estimation of the load over a certain primary substation. Thus, the bottom-up load is directly related to the actual load measurement. The bottom-up is traditionally used by distribution system operators to get an estimate of load on the grid on places where no measurements are available. This bottom-up estimate tries to reconstruct the total load based on telemetry measurements from bulk consumers, from aggregated smaller scale measurements, and from average profiles based on smart meters at consumers’ homes and some smaller bulk consumers. In the latter machine learning is used to estimate the load profiles of those consumers that do not have a smart meter, or have not consented to have their smart meter data read Heres et al. (2023). In order to acquire the final bottom-up load, Alliander uses the SunDance algorithm Chen and Irwin (2017) for disaggregation of net consumption and generation, $k$-means clustering MacQueen et al. (1967); Lloyd (1982) for generating load profiles, and XGBoost Chen and Guestrin (2016) for determining which clustered load profiles should be used instead of missing smart meter measurements Heres et al. (2023). The bottom-up load, in contrast to the actual load, can always be measured in the negative due to being based on more recent, lower-level, measuring equipment. We will use this property in correcting the sign of the load measurement. More details regarding the bottom-up generation methodology can be found in Heres et al. (2023). Most often, the bottom-up load is fairly accurate. Most failure cases are not caused by the algorithm, but rather by incorrect grid-topology data, causing consumers to be wrongfully included or excluded.
@@
In order to discern between a primary substation connection that is net consuming or net producing, the load measurements $S$ are given a sign based on $P$. If a primary substation connection is consuming more than it is producing, the load measurement is assigned a positive sign. A negative sign therefore means that the primary substation is net producing. However, not all primary substations are outfitted with measurement equipment to determine whether the primary substation is net producing or consuming: they just measure the absolute current $I$, thus the sign needs to be corrected later on. As all bottom-up load measurements are based on $P$ measurements and load estimates, they never have a missing sign. A typical primary substation measurement time series, consisting out of the load ($S$), the bottom-up load, and an illustrative example of the needed minimum and maximum capacity is shown in Figure 2, where it should be noted that this measurement does not contain anomalies or switch events. In this figure we can additionally see the minimum and maximum loads ($S$) that are vitally important for the planning of grid expansion. On top of the maximum or minimum load, there should be enough redundant capacity in order to allow for rerouting in case of grid failures. All remaining capacity is unused, and quantification of this unused capacity is essential for determining whether the capacity of a station should be expanded, or whether there is room for additional customers. Should anomalies or switch events be present, we expect to see these minimum and maximum values be inflated or deflated, leading to inaccurate estimates of the unused capacity of a primary substation, thereby interfering with power grid expansion planning. An example of wrong capacity estimates leading to unused capacity can be found in Figure 3
@@
To summarize, at each 15-minute interval $t$ for each station $i$ we measure the load $s_{t}^{i}$ and estimate the bottom-up load $b_{t}^{i}$. We do this for a full year of measurements, yielding for each station $i$ a load vector $\boldsymbol{\mathbf{s}}^{i}\in S$, and a bottom-up load vector $\boldsymbol{\mathbf{b}}^{i}\in B$, where $S$ and $B$ respectively now denote the full collections of load and bottom-up vectors as measured for each of the $180$ measured stations. In order to allow for evaluation of automatic segmentation and anomaly detection algorithms, each 15-minute measurement $y_{t}^{i}$ has additionally been labelled, leading to a detailed labelled segmentation of each time series for each station $i$, $\boldsymbol{\mathbf{y}}^{i}\in Y$, which we will treat as our gold standard. Possible values for this label are: 00 (no anomaly or switch event); $1$ (anomaly or switch event); $5$ the labeler is uncertain whether this should be label 00 or $1$. For simplicity, the dependence of vectors $\boldsymbol{\mathbf{s}}$, $\boldsymbol{\mathbf{b}}$ and $\boldsymbol{\mathbf{y}}$ on index of the station $i$ will be omitted throughout the remainder of this paper, and will only be mentioned explicitly when necessary.
The time series of all primary substations show great variation, both within a primary substation, but also between primary substations. The average load for a primary substation can range from 100’s to 10.000’s of kilowatts. In addition, not all bottom-up loads are equally accurate for each primary substation. This variability leads to a challenging segmentation task.
@@ 2.1.1 Event-length Categories
We use a variety of measures to judge the quality of the time series segmentation, see Section 2.4.1. Generally, these measures are calculated over all individual time points. This can be done when the segments are of somewhat similar length. In this use case however, anomalies are very short events, while switch events might be very long. This disparity in event length is illustrated in Figure 4. From this figure, we can clearly observe that long events are rare, but make up the majority of measured data labelled as $1$, while short events/anomalies are much more frequent, but only make up a small part of the label $1$ data. In order to alleviate this problem while optimizing and evaluating our methods, we divide the anomaly/switch events in 4 categories based on their length based on how many samples they consists of, and calculate all measures, see Section 2.4.1 for each category. The 4 categories of event lengths are defined as: 1 up to and including 24 samples, 25 up to including 288 samples, 289 up to and including 4032 samples, and 4033 samples and longer. From here on out we will refer to these, for sake of clarity, as their equivalents in time units: 15 minutes to 6 hours, 6 hours to 3 days, 3 to 42 days, and 42 days or longer.
@@ 2.2 Preprocessing
In order to make each time series suitable for further analysis, several preprocessing steps have been conducted. First, missing measurements or bottom-up loads are removed from each time series. Missing data can occur when there is a communication error in the system. Communication errors are typically characterized by repeated measurements or bottom-up loads. To correct for the discrepancies between the load measurement and the bottom-up load, we perform a linear regression to better match the two. The linear regression is done on a subset of the time series, specifically by excluding everything outside user-defined quantiles $q_{\textrm{lower}}\%$ and $q_{\textrm{upper}}\%$. This is done to ensure that any anomalies or switch events do not steer the linear regression. Lastly, we perform a sign correction on $S$ measurements as some measurement equipment cannot measure load signs. We list the preprocessing steps as pseudocode in Algorithm 1 in the order in which they are applied to each individual station measurement, i.e., all $\boldsymbol{\mathbf{s}}\in S$ and $\boldsymbol{\mathbf{b}}\in B$. Note that $\boldsymbol{\mathbf{s}}^{i}$ and $\boldsymbol{\mathbf{b}}^{i}$ are vectors of equal length.
@@
6:     if $\textrm{bottomUpMissing}(s_{i})~{}\textbf{or}~{}\textrm{repeatedMeasurements}(s_{i},r)$ then
7:         Remove $s_{i}$ from $\boldsymbol{\mathbf{s}}$ and $b_{i}$ from $\boldsymbol{\mathbf{b}}$
@@
As one can note, this procedure has several user-defined hyperparameters, which can be optimized. These are specifically the range of the quantiles, $q_{\textrm{lower}}\%$ and $q_{\textrm{upper}}\%$, used in the filtering procedure, as well as the number of $\boldsymbol{\mathbf{s}}$ measurements, $r$, that have to be identical in order to be classified as missing due to a communication error. Based on manual observations of the preprocessing procedure, we’ve selected $q_{\textrm{lower}}\%$ and $q_{\textrm{upper}}\%$ to be $10\%$ and $90\%$ respectively, and $r$ to be $5$.
@@ 2.3 Algorithms and Optimization
We specifically make use of unsupervised anomaly detection algorithms. Unsupervised algorithms do not learn from labelled data, but only consider the measurements. We have annotated a fairly large number of $180$ yearly primary substation measurements, but recognize that anomalies and switch events are very rare and heterogeneous events. Because of this, we applied unsupervised methods of detection. We then assume that some higher level hyperparameters of the unsupervised procedure, specifically the thresholds used for acquiring labels, will generalize towards unobserved types of anomalies and switch events, we will evaluate this assumption on the test set.

### v3

# Acquiring Better Load Estimates by Combining Anomaly and Change Point Detection in Power Grid Time Series Measurements

@@ 1 Introduction
Next to expansion, smarter use of the existing grid is essential. This can be done through for example flexible energy contracts, the use of redundant grid capacity for power generation, and the use of batteries (Sufyan et al., 2019). Better insight into grid usage over time is needed to facilitate these changes. With more measurements this can be achieved, but this means also more data that needs to be cleaned before it can be used for analysis.
In order to determine key points of expansion, an accurate overview of the current state of grid capacity needs to be made. In this study, we specifically study primary substation-level measurements on the Dutch power grid managed by Alliander. In order to know what percentage of a primary substation’s capacity is being used, accurate estimates of the minimum and maximum load of the subgrid that substation supplies must be made. This process is more commonly known as demand modeling (Betge et al., 2021). Primary substation-level time series load measurements, however, cannot be directly interpreted without giving a distorted view. These measurements are contaminated with anomalies, for example caused by measurement errors, and with switch events. In a switch event power from different primary substations is rerouted to or from part of the measured primary substation’s route. An example of this can be found in Figure 1, where due to a cable failure power is rerouted from primary substation 1 to supply secondary substations E and F. Rerouting power reduces the blackout time in case of damage to a cable in the power grid, and is possible when the grid is sufficiently redundant to circumvent the broken cable. Switch events can also happen when long-term maintenance is performed on a primary substation or any its cables or secondary substations. Switch events can have a large range of possible lengths, from a few minutes, to multiple months, depending on how fast the underlying problem is resolved. Both of these, measurement errors, which we will call anomalies, and switch events, need to be filtered out to get an accurate estimate of the load profile. This more accurate load profile can be used directly for more optimized usage of the grid. It can then be used to find the true minimum and maximum load of the substation under normal operating conditions, so the redundancy capacity can be added separately. The task of detecting which parts of a load measurement represent the normal situation can be seen as a time series segmentation task.
@@
Traditionally, these load measurement are manually segmented and annotated by domain experts within the Alliander organization. This procedure is however extremely time consuming, and produces annotations with a high percentage of label noise. This label noise is of little concern in determining the minimum and maximum loads, as the annotation procedure is generally optimized in such a way that these loads are as accurate as possible. The resulting annotation can however not be used for other purposes such as year-round usage insights. An automated annotation procedure, capable of producing high-quality annotations, can therefore both free up valuable domain expert time, as well as allow for detailed studies on load measurement data previously unfeasible. Because economically and logistically vital decisions are made based on these load estimates, it is extremely important that any machine learning model automating this procedure is highly interpretable.
For smart grids, change point and anomaly detection algorithms have been broadly applied. Zhang et al. (2021) describe multiple time series anomaly detection algorithms used in various applications across smart grids. They differentiate between point, contextual, and pattern anomalies. Thomas et al. (2020) describe a novel method tailored to detecting islanding. In addition to islanding, their method, which uses $k$-means clustering and empirical mode decomposition, is able to detect load switch events and other faults. Neagoe et al. (2021) used change point detection in order to detect multi-year patterns in hydropower generation in Romania. Wang and Ahn (2020) combine a regression-based anomaly detection model with SVM, $k$NN and a cross-entropy loss function to detect anomalies in a Tanzanian solar power plant. Wang et al. (2023) present methodology for simultaneously performing load forecasting and semi-supervised anomaly detection and apply it to the UCI electric load dataset. They report higher specificity and sensitivity than neural network-based methods, but find similar F1 scores. For shorter time series, Rajabi et al. (2020) have compared various clustering methods for load pattern segmentation.
@@ 2.1 Data
In this study we optimize and evaluate our algorithms for anomaly and switch event detection on a total of $180$ primary substation load measurements. Most primary substation measurements span a full year in length, and provide measurements at regular 15-minute intervals of the apparent power $S$, which we will refer to as load. We calculate $S$ from the active power $P$ and the reactive power $Q$ and assign it the sign of $P$, $S=sign(P)*\sqrt{P^{2}+Q^{2}}$. In some cases, the measurement equipment of a primary substation does not allow for accurate measurements of both $P$ and $Q$. In these cases, $S$ is calculated from $S=\sqrt{3}*V*I$. The $\sqrt{3}$ term originates from the fact that in a 3-phase system the phase voltage rather than the line voltage is used.
@@
At each 15-minute interval over which the load is measured, we also calculate the so-called bottom-up load $B$ throughout the subgrid supplied by that specific primary substation. The bottom-up load is an estimation of the load over a certain primary substation, but not a direct measurement, like $S$. Thus, the bottom-up load is directly related to the actual load measurement. The bottom-up is traditionally used by distribution system operators to get an estimate of load on the grid on places where no measurements are available. This bottom-up estimate tries to reconstruct the total load based on telemetry measurements from bulk consumers, from aggregated smaller scale measurements, and from average profiles based on smart meters at consumers’ homes and some smaller bulk consumers. In the latter machine learning is used to estimate the load profiles of those consumers that do not have a smart meter, or have not consented to have their smart meter data read (Heres et al., 2023). In order to acquire the final bottom-up load, Alliander uses the SunDance algorithm (Chen and Irwin, 2017) for disaggregation of net consumption and generation, $k$-means clustering (MacQueen et al., 1967; Lloyd, 1982) for generating load profiles, and XGBoost (Chen and Guestrin, 2016) for determining which clustered load profiles should be used instead of missing smart meter measurements (Heres et al., 2023). More details regarding the bottom-up generation methodology can be found in (Heres et al., 2023). Most often, the bottom-up load is fairly accurate. Most failure cases are not caused by the algorithm, but rather by incorrect grid-topology data, causing consumers to be wrongfully included or excluded.
@@
In order to discern between a primary substation connection that is net consuming or net producing, the load measurements $S$ are given a sign based on $P$. If a primary substation connection is consuming more than it is producing, the load measurement is assigned a positive sign. A negative sign therefore means that the primary substation is net producing. However, not all primary substations are outfitted with measurement equipment to determine whether the primary substation is net producing or consuming: they just measure the absolute current $I$, thus the sign needs to be corrected later on. The bottom-up load, in contrast to the actual load, can always be measured in the negative due to being based on more recent, lower-level, measuring equipment. We will use this property in correcting the sign of the load measurement. Bottom-up load measurements are furthermore based on $P$ measurements and load estimates so they never have a missing sign. A typical primary substation measurement time series, consisting out of the load ($S$), the bottom-up load, and an illustrative example of the needed minimum and maximum capacity is shown in Figure 2, where it should be noted that this measurement does not contain anomalies or switch events. In this figure we can additionally see the minimum and maximum loads ($S$) that are vitally important for the planning of grid expansion. On top of the maximum or minimum load, there should be enough redundant capacity in order to allow for rerouting in case of grid failures. All remaining capacity is unused, and quantification of this unused capacity is essential for determining whether the capacity of a station should be expanded, or whether there is room for additional customers. Should anomalies or switch events be present, we expect to see these minimum and maximum values be inflated or deflated, leading to inaccurate estimates of the unused capacity of a primary substation, thereby interfering with power grid expansion planning. An example of wrong capacity estimates leading to unused capacity can be found in Figure 3.
@@
To summarize, at each 15-minute interval $t$ for each station $i$ we measure the load $s_{t}^{i}$ and estimate the bottom-up load $b_{t}^{i}$. We do this for a full year of measurements, yielding for each station $i$ a load vector $\boldsymbol{\mathbf{s}}^{i}\in S$, and a bottom-up load vector $\boldsymbol{\mathbf{b}}^{i}\in B$, where $S$ and $B$ respectively now denote the full collections of load and bottom-up vectors as measured for each of the $180$ measured stations. In order to allow for evaluation of automatic segmentation and anomaly detection algorithms, each 15-minute measurement $y_{t}^{i}$ has additionally been labeled, leading to a detailed labeled segmentation of each time series for each station $i$, $\boldsymbol{\mathbf{y}}^{i}\in Y$, which we will treat as our gold standard. Possible values for this label are: 00 (no anomaly or switch event); $1$ (anomaly or switch event); $5$ (the labeler is uncertain whether this should be label 00 or $1$). For simplicity, the dependence of vectors $\boldsymbol{\mathbf{s}}$, $\boldsymbol{\mathbf{b}}$ and $\boldsymbol{\mathbf{y}}$ on index of the station $i$ will be omitted throughout the remainder of this paper, and will only be mentioned explicitly when necessary.
The time series of all primary substations show great variation, both within a primary substation, but also between primary substations. The average load for a primary substation can range from 100’s to 10,000’s of kilowatts. In addition, not all bottom-up loads are equally accurate for each primary substation. This variability leads to a challenging segmentation task.
@@ 2.1.1 Event-length Categories
We use a variety of measures to judge the quality of the time series segmentation, see Section 2.4.1. Generally, these measures are calculated over all individual time points. This can be done when the segments are of somewhat similar length. In this use case however, anomalies are very short events, while switch events might be very long. This disparity in event length is illustrated in Figure 4. From this figure, we can clearly observe that long events are rare, but make up the majority of measured data labeled as $1$, while short events/anomalies are much more frequent, but only make up a small part of the label $1$ data. In order to alleviate this problem while optimizing and evaluating our methods, we divide the anomaly/switch events in 4 categories based on their length based on how many samples they consists of, and calculate all measures, see Section 2.4.1 for each category. The 4 categories of event lengths are defined as: 1 up to and including 24 samples, 25 up to including 288 samples, 289 up to and including 4032 samples, and 4033 samples and longer. From here on out we will refer to these, for sake of clarity, as their equivalents in time units: 15 minutes to 6 hours, 6 hours to 3 days, 3 to 42 days, and 42 days or longer. Roughly speaking the categories contain the following types of events: 15 minutes to 6 hours generally contains measurement errors, 6 hours to 3 days contains longer measurement errors and very short switch events that are easily resolved, 3 to 42 days contains short switch events, as well as more complex rerouting, and 42 days or longer contains switch events caused by for example complex rerouting, long-term maintenance, and grid expansion. It should be noted that this categorization is not set in stone, and may be chosen slightly differently depending on the prevalence of event types in other applications.
@@ 2.2 Preprocessing
In order to make each time series suitable for further analysis, several preprocessing steps have been conducted. First, missing $S$ measurements and corresponding bottom-up loads are removed from each time series. Missing data can occur when there is a communication error in the system. Communication errors are typically characterized by repeated measurements or bottom-up loads. To correct for the discrepancies between the load measurement and the bottom-up load, we perform a linear regression to better match the two. The multiplication term corrects for multiplicative mismatches caused by over- or underestimating the amount of load of the customers and the grid losses that depend on the load (copper-losses), while the additive baseline correction corrects for constant grid losses, or iron losses, and, in the case of a current measurement, for the constant reactive power caused mainly by the capacitance of the cables. The linear regression is done on a subset of the time series, specifically by excluding everything outside user-defined quantiles $q_{\textrm{lower}}\%$ and $q_{\textrm{upper}}\%$. This is done to prevent that any anomalies or switch events steer the linear regression. Lastly, we perform a sign correction on $S$ measurements as some measurement equipment cannot measure load signs. We list the preprocessing steps as pseudocode in Algorithm 1 in the order in which they are applied to each individual station measurement, i.e., all $\boldsymbol{\mathbf{s}}\in S$ and $\boldsymbol{\mathbf{b}}\in B$. Note that $\boldsymbol{\mathbf{s}}^{i}$ and $\boldsymbol{\mathbf{b}}^{i}$ are vectors of equal length. The used functions “bottomUpMissing” and “repeatedMeasurements” are described in Algorithm 2 and Algorithm 3 respectively.
@@
6:     if $\textrm{bottomUpMissing}(b_{i})~{}\textbf{or}~{}\textrm{repeatedMeasurements}(s,i,r)$ then
7:         Remove $s_{i}$ from $\boldsymbol{\mathbf{s}}$ and $b_{i}$ from $\boldsymbol{\mathbf{b}}$triangleright See Algorithm 2 and 3.
@@
Algorithm 2 bottomUpMissing
1:Input: A single bottom-up measurement $b_{i}$
2:Output: Boolean $\beta$ indicating whether the bottom-up load measurement $b_{i}$ is missing
3:
4:
$\beta\leftarrow\begin{cases}\textrm{True},&\text{if }b_{i}=\text{NaN}\\ \textrm{False},&\text{otherwise}\end{cases}$
Algorithm 3 repeatedMeasurements
1:Input: Measurement $\boldsymbol{\mathbf{s}}$ and index $i$
2:Hyperparameters: The maximum number of repeated measurements $r$
3:Output: Boolean $\beta$ indicating whether the maximum allowed number of repeated measurements and adjacent to $s_{i}$ is exceeded
4:
5:$\beta\leftarrow\textrm{False}$
6:$c\leftarrow 0$triangleright Repeated measurement counter
[…]
8:for $j\leftarrow\max(i-r,0)$ to $\min(i+r,n)$ do $\triangleright$ $n$ is the length of the vector $\boldsymbol{\mathbf{s}}$
9:     if $s_{j}=s_{i}$ then
10:         $(c,\textit{max\_c})max_$_c )
11:     else
12:         $c\leftarrow 0$
13:     end if
14:end for
[…]
16:     $\beta\leftarrow\textrm{True}$
17:end if
As one can note, this procedure has several user-defined hyperparameters, which can be optimized. These are specifically the range of the quantiles, $q_{\textrm{lower}}\%$ and $q_{\textrm{upper}}\%$, used in the filtering procedure, as well as the number of $\boldsymbol{\mathbf{s}}$ measurements, $r$, that have to be identical in order to be classified as missing due to a communication error. Based on manual observations of the preprocessing procedure, we have selected $q_{\textrm{lower}}\%$ and $q_{\textrm{upper}}\%$ to be $10\%$ and $90\%$ respectively, and $r$ to be $5$.
@@ 2.3 Algorithms and Optimization
We specifically make use of unsupervised anomaly detection algorithms. Unsupervised algorithms do not learn from labeled data, but only consider the measurements. We have annotated a fairly large number of $180$ yearly primary substation measurements, but recognize that anomalies and switch events are very rare and heterogeneous events. Because of this, we applied unsupervised methods of detection. We then assume that some higher level hyperparameters of the unsupervised procedure, specifically the thresholds used for acquiring labels, will generalize towards unobserved types of anomalies and switch events, we will evaluate this assumption on the test set.